# Leveraging Offline Data in Linear Latent Bandits

### v1

@@ Abstract
Sequential decision-making domains such as recommender systems, healthcare and education often have unobserved heterogeneity in the population that can be modeled using latent bandits – a framework where an unobserved latent state determines the model for a trajectory. While the latent bandit framework is compelling, the extent of its generality is unclear. We first address this by establishing a de Finetti theorem for decision processes, and show that every exchangeable and coherent stateless decision process is a latent bandit. The latent bandit framework lends itself particularly well to online learning with offline datasets, a problem of growing interest in sequential decision-making. One can leverage offline latent bandit data to learn a complex model for each latent state, so that an agent can simply learn the latent state online to act optimally. We focus on a linear model for a latent bandit with $d_{A}$-dimensional actions, where the latent states lie in an unknown $d_{K}$-dimensional subspace for $d_{K}\ll d_{A}$. We present SOLD, a novel principled method to learn this subspace from short offline trajectories with guarantees. We then provide two methods to leverage this subspace online: LOCAL-UCB and ProBALL-UCB. We demonstrate that LOCAL-UCB enjoys $\tilde{O}(\min(d_{A}\sqrt{T},d_{K}\sqrt{T}(1+\sqrt{d_{A}T/d_{K}N})))$ regret guarantees, where the effective dimension is lower when the size $N$ of the offline dataset is larger. ProBALL-UCB enjoys a slightly weaker guarantee, but is more practical and computationally efficient. Finally, we establish the efficacy of our methods using experiments on both synthetic data and real-life movie recommendation data from MovieLens.
@@ 1 Introduction
Many sequential-decision making problems can be effectively modeled using the bandit framework. This can span domains as diverse as healthcare (Lu et al., 2021), randomized clinical trials (Press, 2009), search and recommendation (Li et al., 2010), distributed networks (Kar et al., 2011), and portfolio design (Brochu et al., 2011). However, in many applications, there often exists unobserved heterogenity in the population, or unobserved context, that influences the distribution of rewards. In Hong et al. (2020), it is shown that this uncertainty can be modeled by a latent bandit (or mixture of bandits). This is a bandit where an unobserved latent state determines the reward model for the trajectory. The latent bandit framework therefore has high practical value.
However, the extent of the generality of this framework is unclear. Given a general decision process, how strong do the conditions need to be to make it a latent bandit? To answer this, we turn to de Finetti’s celebrated theorem, which says that exchangeable sequences of Bernoulli random variables are mixtures of sequences of i.i.d. Bernoulli random variables. This has also been extended to larger classes of stochastic processes. For example, Diaconis and Freedman (1980) show that any random process satisfying recurrence and a weakened notion of exchangeability is a mixture of Markov chains. Inspired by such work, we define a general stateless decision process, and consider a weak notion of exchangeability as well as a natural coherence condition. We show that in the spirit of de Finetti’s theorem, any coherent and exchangeable stateless decision process is a latent bandit.
In light of this equivalence and the empirical relevance of linear contextual bandits (Li et al., 2010; Abbasi-Yadkori et al., 2011; Abbasi-yadkori et al., 2011; Abbasi-Yadkori et al., 2012), we define and explore the setting of linear latent contextual bandits, where a low-dimensional latent state imposes a linear effect on the reward model parameters. When interacting with a latent bandit online, an agent receives rewards based on a single latent state. For example, a patient’s underlying genetic makeup could be the latent parameter that affects their response to treatment. Often, one also has access to large offline datasets with unobserved heterogeneity, like data from randomized controlled trials. Using traditional bandit algorithms in this setting does not leverage the offline data available to the agent. One also cannot simply treat the offline data as coming from a single bandit, since there is significant distribution shift between the mean offline parameter and a single online parameter.
Addressing this, Hong et al. (2020) considers the problem of speeding up online learning using models learnt offline for finitely many latent states. Unlike them, we tackle the problem of using subspaces learnt offline to accelerate online learning, as learning the "latent subspace" spanned by possible reward model parameters is equivalent to learning the infinitely many models in the subspace. We provide a method to learn this subspace from a dataset of trajectories collected under some behavior policy $\pi_{b}$, and seek to use it to speed up online learning. Our subspace learning algorithm is inspired by a long line of work in using spectral methods to learn mixtures of various generative models (Vempala and Wang, 2004; Kong et al., 2020; Chen and Poor, 2022; Kausik et al., 2023). However, each of these methods deals with finitely many latent states, and clusters data under separation assumptions. Clustering is not compatible with continuous latent states, and so, novel ideas are needed to extend these spectral methods to our setting. Finally, the subspace learnt offline is typically an uncertain estimate of the true subspace of reward parameters. Online methods using the subspace should thus account for this uncertainty. We therefore design two such methods with a tractability-tightness tradeoff. With this in mind, we outline our contributions below:
  • We are the first, to our knowledge, to prove a de Finetti theorem for decision processes.
@@
  • We present SOLD, an offline method for learning low-dimensional subspaces of reward parameters with guarantees, generalizing spectral learning methods from discrete to continuous latent states.
@@
  • We present LOCAL-UCB, an online algorithm leveraging subspaces and uncertainty sets estimated offline to sharpen optimism for LinUCB, with $\tilde{O}(\min(d_{A}\sqrt{T},d_{K}\sqrt{T}(1+\sqrt{d_{A}T/d_{K}N})))$ regret.
@@
  • Finally, we present ProBALL-UCB, a practical and computationally efficient online algorithm with a slightly looser regret guarantee. This illustrates a general algorithmic idea for integrating offline subspace estimation into optimistic algorithms.
@@
  • We establish the efficacy of our algorithms outlined above through a simulation study and a demonstration on a real recommendation problem with the MovieLens-1M (Harper and Konstan, 2015) dataset.
Other related work.
The work of Tennenholtz et al. (2021) considers linear contextual bandits with partially observed offline data and fully observable online features, albeit in a very different model of partial information. Additionally, there has been an increasingly large amount of work in offline-to-online transfer learning for reinforcement learning, also known as hybrid RL (Song et al., 2023; Nakamoto et al., 2023; Amortila et al., 2024; Li et al., 2023; Rajeswaran et al., 2017; Hester et al., 2018; Nair et al., 2018, 2020; Ball et al., 2023; Xie et al., 2022; Ren et al., 2024; Wagenmaker and Pacchiano, 2023; Tan and Xu, 2024).
2 A De Finetti Theorem for Decision Processes
We first define a stateless decision process at a high level of generality. [^1] Liu et al. (2023) work with a much more restrictive notion of a generalized bandit and use the original de Finetti theorem in some of their lemmas. See Appendix B.1 for a discussion.
@@ 2 A De Finetti Theorem for Decision Processes
A stateless decision process (SDP) with action set $\mathcal{A}$ is a probability space $(\Omega,\mathcal{G},\mathbb{P})$ equipped with a family of random maps $\mathcal{F}_{H}:\Omega\to(\mathcal{A}^{H}\to\mathbb{R}^{H})$ for $H\in\mathbb{N}\cup\{\infty\}$.
That is, given a sequence of actions $(a_{1},\dots a_{H})$, an SDP generates a random sequence of rewards $(Y_{1},\dots Y_{H})$. In the light of this interpretation, we will abuse notation to denote by $\mathcal{F}_{H}(a_{1},\dots a_{H})$ the random variable $\omega\mapsto\mathcal{F}_{H}(\omega)(a_{1},\dots a_{H})$. Without any coherence between $\mathcal{F}_{H}$ across $H$, a stateless process can behave arbitrarily for different horizons $H$. We present a natural coherence condition below, essentially requiring that a given action sequence should produce consistent rewards.
Definition 2.
A stateless decision process is said to be coherent if for any $h\leq k\leq H,H^{\prime}\in\mathbb{N}\cup\{\infty\}$ and for any two action sequences $\tau,\tau^{\prime}$ of lengths $H$ and $H^{\prime}$ sharing the same actions $(a_{h},\dots a_{k})$ from index $h$ to $k$, with $\mathcal{F}_{H}(\tau)=(Y_{1},\dots Y_{H})$ and $\mathcal{F}_{H^{\prime}}(\tau^{\prime})=(Y_{1}^{\prime},\dots Y_{H^{\prime}}^{\prime})$, we have $(Y_{h},\dots Y_{k})=(Y_{h}^{\prime},\dots Y_{k}^{\prime})$, viewed as functions of $\Omega$.
It is natural to require equality in value and not just in distribution, since after taking an extra action, the values of past rewards stay the same, not just their distribution. We also give a natural definition for exchangeability of a stateless decision process – namely that exchanging any two rewards should lead to the distribution obtained by exchanging the corresponding actions.
Definition 3.
A stateless decision process is said to be exchangeable if for any permutation $\pi:[H]\to[H]$ and $\mathcal{F}_{H}(a_{1},\dots a_{H})=(Y_{1},\dots Y_{H})$, we have $\mathcal{F}_{h}(a_{\pi(1)},\dots a_{\pi(H)})\sim(Y_{\pi(1)},\dots Y_{\pi(H)})$.
Definition 4.
A latent bandit is a stateless decision process equipped with a random measure-valued function $F:\Omega\to(\mathcal{A}\to\mathcal{P}(\mathbb{R}))$ so that for any $H$ and action sequence $(a_{1},\dots a_{H})$, the rewards $(Y_{1},\dots Y_{H}):=\mathcal{F}_{H}(a_{1},\dots a_{H})$ are independent conditioned on $F$. Moreover, the conditional distribution $\mathcal{L}[Y_{h}\mid F]=F(a_{h})$ for all $h\leq H$ [^2] We abuse notation twice here. First, we write $F(a_{h}):=(\omega\mapsto F(\omega)(a_{h}))$. Second, as the regular conditional distribution $\mathcal{L}[Y_{h}\mid F]$ is a kernel that maps from $\Omega\times\mathcal{B}\to\mathbb{R}$, we view $F(a_{h})$ as its curried map $(\omega,B)\mapsto F(a_{h})(\omega)(B)$. A discussion of issues like measurability and well-definedness is in Appendix B.2.
While exchangeability and coherence are reasonable conditions on an SDP and are clearly satisfied by latent bandits, it is a-priori unclear if they are sufficient to ensure that the SDP is a latent bandit. In particular, since only exchanging rewards from the same action preserves the distribution, standard de Finetti proof ideas do not immediately apply. After all, it is possible that an SDP could be cleverly designed to choose rewards adaptively across time and satisfy these properties. Reassuringly, no such counterexamples exist and we have the following theorem.
Theorem 1 (De Finetti Theorem for Stateless Decision Processes).
We show in Lemma 2 that coherence is not a consequence of exchangeability – it is a necessary condition for being a latent bandit. Finally, we consider contexts and define contextual decision processes in a manner agnostic to context transitions. That is, the process only carries the data of how rewards are generated from given sequences of contexts and actions, while the context transitions themselves may be generated by a different process.
Definition 5.
It is natural to isolate away context dynamics for processes like stochastic contextual bandits, in which actions do not affect the context transitions – in contrast to transition-aware definitions in Jiang et al. (2016). We can analogously define coherence and exchangeability for TACDPs and define latent contextual bandits by simply replacing $\mathcal{A}$ with $\mathcal{X}\times\mathcal{A}$ in the definitions above. We then show an analogous de Finetti theorem, as a corollary of our proof of Theorem 1. See Appendix B.2 for more details.
3 Linear Latent Contextual Bandits
Latent bandits are a powerful and general framework for encoding uncertainty in reward models. However, this generality is both a blessing and a curse. Without further assumptions on the structure of the latent state $F$, one cannot hope to estimate it. We therefore focus on a special case here. We consider the natural generalization of the linear contextual bandit to the latent bandit setting, where we impose a linear structure on the effect of latent states. We further justify this by noting that in most application domains, it is reasonable to assume that a parsimonious, low-dimensional latent state affects the reward distribution. This motivates the following definition.
Definition 6.
@@ 3 Linear Latent Contextual Bandits
A linear latent contextual bandit is a TACDP equipped with a feature map for context-action pairs $\phi:\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d_{A}}$, a latent random variable $\bm{\theta}\in\mathbb{R}^{d_{K}}$ with distribution $\mathcal{D}_{\theta}$ and a map $\textbf{U}_{\star}:\mathbb{R}^{d_{K}}\to\mathbb{R}^{d_{A}}$ such that for any $H$ and context-action sequence $((x_{1},a_{1}),\dots(x_{H},a_{H}))$, the rewards $(Y_{1},\dots Y_{H})$ are independent conditioned on $\bm{\theta}$. Moreover, $Y_{h}\mid\bm{\theta}\sim\phi(x_{h},a_{h})^{\top}\bm{\beta}+\epsilon$, where $\epsilon$ is subgaussian noise independent of all actions and all other observations, and $\bm{\beta}=\textbf{U}_{\star}\bm{\theta}$.
@@
In other words, this is a latent contextual bandit where the random measure-valued function $F:\Omega\to((\mathcal{X}\times\mathcal{A})\to\mathcal{P}(\mathbb{R}))$ is defined by setting $F(\omega)(x,a)$ to be the distribution given by $\phi(x,a)^{\top}\textbf{U}_{\star}\bm{\theta}(\omega)+\epsilon$, for any $\omega\in\Omega$ and $(x,a)\in(\mathcal{X}\times\mathcal{A})$. Further, we note that WLOG $\textbf{U}_{\star}$ is unitary. This is because for any invertible map $A:\mathbb{R}^{d_{K}}\to\mathbb{R}^{d_{K}}$, the observation distribution does not change upon replacing $\bm{\theta}$ with $A\bm{\theta}$ and replacing $\textbf{U}_{\star}$ with $\textbf{U}_{\star}A^{-1}$. One can see this as a generalization of the fact that with finitely many latent states, permuting the labels and rewards together does not change observations.
@@
Let us now assume that we have access to a dataset $\mathcal{D}_{\operatorname{off}}$ of $N$ short trajectories $\tau_{n}=((x_{n,1},a_{n,1},r_{n,1}),...,$ $(x_{n,H},a_{n,H},r_{n,1}))$ of length $H$, collected by some behavior policy $\pi_{b}$. In online deployment, a single latent label $\bm{\theta}_{\star}$ is chosen and rewards are generated using $\bm{\beta}_{\star}=\textbf{U}_{\star}\bm{\theta}_{\star}$. At each timestep $t$, an agent observes contexts $x_{t}$ and uses both the offline data and the online data at time $t$ to execute a policy $\pi_{t}$. Define the optimal action at time $t$ by $a^{\star}_{t}:=\max_{a}\phi(x_{t},a_{t})^{\top}\bm{\beta}_{\star}$ We tackle the problem of minimizing the frequentist regret in linear latent contextual bandits, given by
@@
Challenges with latent bandits. Despite the linear assumption, and the dimension reduction obtained in the common case when $d_{K}\ll d_{A}$, significant challenges remain. First, the value of the latent state $\bm{\theta}$ and the map $\textbf{U}_{\star}$ are both unknown a priori, making it hard to leverage the low-dimensional structure of the problem. Second, a good choice of dimension $d_{K}$ is itself unknown a priori, and must be determined from data in a principled manner. In the following section, we will provide a method to estimate latent subspaces from offline data that allows us to overcome these challenges.
4 Estimating Latent Subspaces from Offline Data

### v2

# Leveraging Offline Data in Linear Latent Contextual Bandits

@@ Abstract
Leveraging offline data is an attractive way to accelerate online sequential decision-making. However, it is crucial to account for latent states in users or environments in the offline data, and latent bandits form a compelling model for doing so. In this light, we design end-to-end latent bandit algorithms capable of handing uncountably many latent states. We focus on a linear latent contextual bandit – a linear bandit where each user has its own high-dimensional reward parameter in $\mathbb{R}^{d_{A}}$, but reward parameters across users lie in a low-rank latent subspace of dimension $d_{K}\ll d_{A}$. First, we provide an offline algorithm to learn this subspace with provable guarantees. We then present two online algorithms that utilize the output of this offline algorithm to accelerate online learning. The first enjoys $\tilde{O}(\min(d_{A}\sqrt{T},d_{K}\sqrt{T}(1+\sqrt{d_{A}T/d_{K}N})))$ regret guarantees, so that the effective dimension is lower when the size $N$ of the offline dataset is larger. We prove a matching lower bound on regret, showing that our algorithm is minimax optimal. The second is a practical algorithm that enjoys only a slightly weaker guarantee, but is computationally efficient. We also establish the efficacy of our methods using experiments on both synthetic data and real-life movie recommendation data from MovieLens. Finally, we theoretically establish the generality of the latent bandit model by proving a de Finetti theorem for stateless decision processes.
@@ 1 Introduction
Many sequential-decision making problems can be effectively modeled using the bandit framework. This can span domains as diverse as healthcare Lu et al., 2021b , randomized clinical trials Press, (2009), search and recommendation Li et al., (2010), distributed networks Kar et al., (2011), and portfolio design Brochu et al., (2011). There is often a wealth of offline data in such domains, which has led to a growing interest in using offline data to accelerate online learning. However, there often also exist unobserved contexts in the population that influence the distribution of rewards, making it non-trivial to leverage offline data. In Hong et al., (2020), it is shown that this uncertainty can be modeled by a latent bandit (or mixture of bandits). This is a bandit where an unobserved latent state determines the reward model for the trajectory. For example, a patient’s underlying genetic conditions in healthcare and a user’s tastes in recommendation systems are both examples of latent states in sequential decision making. Typically, these latent states are less complex than the actual models underlying users or patients, making it valuable to reduce the online task to learning the latent state Hong et al., (2020, 2022).
The latent bandit framework therefore has high practical value, and efficient principled algorithms are needed for using offline data to speed up online learning. Using traditional bandit algorithms in this setting does not leverage the offline data that is often available to the agent. Naturally, one also cannot treat the offline data as coming from a single bandit. For example, different user tastes or different underlying genetic conditions require modeling the offline data as coming from a latent bandit. So, we have to develop algorithms specific to the latent bandit setting that leverage the offline data to improve online performance.
We note that in bandit literature, it is common to impose a structure on bandit rewards when designing algorithms, the most popular one being a linear structure Li et al., (2010); Abbasi-Yadkori et al., (2011). In this light, we study a linear contextual bandit setting where each user has its own high-dimensional reward parameter, but reward parameters across users lie in a low-rank subspace. This is a linear latent contextual bandit [^1] This can also be thought of as a continuous mixture of bandits., and is much more general than existing models that restrict themselves to finitely many latent states Hong et al., (2020, 2022). We design a two-pronged algorithm to tackle this setting. First, we provide a method to approximate the low-dimensional subspace spanned by latent states from an offline dataset of unlabeled trajectories collected under some behavior policy $\pi_{b}$. This is non-trivial since the trajectories are unlabeled, and standard unsupervised learning methods fail. Second, we use this subspace to speed up online learning. However, since the subspace is only learnt approximately, we also tackle the non-trivial task of accounting for the uncertainty in the subspace. We design two methods for the latter, facing a trade-off between computational tractability and tightness of guarantees. Experiments show the efficacy of our methods.
While latent bandits have thus shown to be a powerful and tractable framework for accounting for uncertainty in reward models, the extent of their generality is unclear. Are there other stateless decision processes that generalize over latent bandits? We end by theoretically demonstrating that under very reasonable assumptions, the answer is no. We show a de Finetti theorem for decision processes, demonstrating that every "coherent" and "exchangeable" stateless (contextual) decision process is a latent (contextual) bandit. With this in mind, we outline our contributions below:
@@
  • Offline method: We present SOLD, a novel offline method for learning low-dimensional subspaces of reward parameters with guarantees, inspired by the novel spectral methods in Kausik et al., (2023).
@@
  • Tight online algorithm: We present LOCAL-UCB, an online algorithm leveraging the subspace estimated offline to sharpen optimism, achieving $\tilde{O}(\min(d_{A}\sqrt{T},d_{K}\sqrt{T}(1+\sqrt{d_{A}T/d_{K}N})))$ regret.
  • Lower Bound: We establish a matching lower bound showing that LOCAL-UCB is minimax optimal. To the best of our knowledge, this is the first lower bound in a hybrid (offline-online) sequential decision-making setting.
@@
  • Tractable online algorithm: Finally, we present ProBALL-UCB, a practical and computationally efficient online algorithm with a slightly looser regret guarantee. This also illustrates a general algorithmic idea for integrating offline subspace estimation into optimistic algorithms.
@@
  • Experiments: We establish the efficacy of our algorithms outlined above through a simulation study and a demonstration on a real recommendation problem with the MovieLens-1M Harper and Konstan, (2015) dataset.
  • Theoretical generality: We are the first, to our knowledge, to prove a de Finetti theorem for decision processes. This establishes the generality of the latent bandit model.
Related work. There are three main threads of related work.
  • Latent Bandits. The line of work most relevant to us has been on latent bandits. The work of Hong et al., (2020, 2022) studies the latent bandit problem under finitely many states. However, they black-box the offline step and do not provide end-to-end guarantees, and their ideas do not extend to infinitely many states. Our work seeks to provide end-to-end guarantees for both the offline and online component under infinitely many latent states.
  • Meta learning, multi-task learning and mixture learning. A long line of work studies learning with multiple underlying tasks or models. For example, the work of Vempala and Wang, (2004); Kong et al., (2020); Anandkumar et al., (2014); Tripuraneni et al., (2022) study learning under latent variable or multi-task models in a supervised setting. The work of Kausik et al., (2023); Chen and Poor, (2022) extend some of these ideas to unsupervised but purely offline learning in a time-series setting. On the other hand, work like Yang et al., (2022); Cella et al., (2022) instead focuses on the purely online setting of learning the low-rank structure while simultaneously interacting with multiple finitely many bandit instances. Finally, Zhou et al., (2024); Lu et al., 2021a work with multiple underlying models in MDPs but in a purely offline and generative model setting respectively. Unlike these papers, we crucially combine the offline and online settings for sequential data and study the problem of using offline data to accelerate online learning in bandits.
  • Hybrid (offline-online) RL. Work in hybrid RL studies the use of offline data to accelerate online RL, first proposed by Song et al., (2023), with extensions to linear MDPs by Wagenmaker and Pacchiano, (2023); Tan et al., (2024). Cai et al., (2024) studies the same problem for contextual bandits. However, all work so far assumes that the offline data is generated by a single model, and does not account for latent states. Our work explores a hybrid offline-online setting while also accounting for the offline data being generated by multiple underlying models.
2 Linear Bandits With Latent Structure
We will first introduce latent contextual bandits, and then specialize to our linear model later in this section. A latent contextual bandit is a decision process with contexts $\mathcal{X}$, actions $\mathcal{A}$ and an random latent state $\theta$ that is sampled independently at the beginning of each trajectory. Given a sequence of contexts and actions, the rewards at all steps are independent conditioned on $\theta$, and depend on the latent state $\theta$, the context and the action. Since we often have access to an offline dataset of trajectories coming from different kinds of users or patients, it is important to account for a changing latent state $\theta$ between trajectories.
As we will see in Section 8, latent bandits are a powerful and general framework for encoding uncertainty in reward models. However, this generality is both a blessing and a curse. It is hard to design concrete algorithms without further assumptions on the structure and effect of the latent state $\theta$. We therefore focus on a linear structure here. We consider the natural generalization of the linear contextual bandit to the latent bandit setting, where we impose a linear structure on the effect of low-dimensional latent states. We further justify this by noting that in most application domains, it is reasonable to assume that a parsimonious, low-dimensional latent state affects the reward distribution. This motivates the following definition.
@@ 2 Linear Bandits With Latent Structure
A linear latent contextual bandit is a linear bandit equipped with a feature map for context-action pairs $\phi:\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d_{A}}$, a latent random variable $\bm{\theta}\in\mathbb{R}^{d_{K}}$ with distribution $\mathcal{D}_{\theta}$ and a map $\textbf{U}_{\star}:\mathbb{R}^{d_{K}}\to\mathbb{R}^{d_{A}}$ such that for any $H$ and context-action sequence $((x_{1},a_{1}),\dots(x_{H},a_{H}))$, the rewards $(Y_{1},\dots Y_{H})$ are independent conditioned on $\bm{\theta}$. Moreover, $Y_{h}\mid\bm{\theta}\sim\phi(x_{h},a_{h})^{\top}\bm{\beta}+\epsilon$, where $\epsilon$ is subgaussian noise independent of all actions and all other observations, and $\bm{\beta}=\textbf{U}_{\star}\bm{\theta}$.
@@
Further, we note that WLOG $\textbf{U}_{\star}$ is unitary. This is because for any invertible map $A:\mathbb{R}^{d_{K}}\to\mathbb{R}^{d_{K}}$, the observation distribution does not change upon replacing $\bm{\theta}$ with $A\bm{\theta}$ and $\textbf{U}_{\star}$ with $\textbf{U}_{\star}A^{-1}$. One can see this as a generalization of the fact that with finitely many latent states, permuting the labels and rewards together does not change observations.
@@
Let us now assume that we have access to a dataset $\mathcal{D}_{\operatorname{off}}$ of $N$ short trajectories $\tau_{n}=((x_{n,1},a_{n,1},r_{n,1}),$ $(x_{n,H},a_{n,H},r_{n,1}))$ of length $H$, collected by some behavior policy $\pi_{b}$. The trajectories are short in the sense that in most relevant domains, individual trajectories are not long enough to learn the underlying reward model. Each trajectory $\tau_{n}$ has a different $\bm{\beta}_{n}=\textbf{U}_{\star}\bm{\theta}_{n}$. In online deployment, a single latent label $\bm{\theta}_{\star}$ is chosen and rewards are generated using $\bm{\beta}_{\star}=\textbf{U}_{\star}\bm{\theta}_{\star}$. At each timestep $t$, an agent observes contexts $x_{t}$ and uses both the offline data and the online data at time $t$ to execute a policy $\pi_{t}$. Define the optimal action at time $t$ by $a^{\star}_{t}:=\max_{a}\phi(x_{t},a_{t})^{\top}\bm{\beta}_{\star}$ We tackle the problem of minimizing the frequentist regret in linear latent contextual bandits, given by
@@
Challenges with latent bandits. Despite the linear assumption, and the dimension reduction obtained in the common case when $d_{K}\ll d_{A}$, significant challenges remain. First, the value of the latent state $\bm{\theta}$ and the map $\textbf{U}_{\star}$ are both unknown a priori, making it hard to leverage the low-dimensional structure of the problem. Second, a good choice of dimension $d_{K}$ is itself unknown a priori, and must be determined from data in a principled manner. Third, even if we learn the low-dimensional structure, our learning will be approximate, and the online procedure must account for this uncertainty. In the following sections, we will provide a method to estimate and use latent subspaces given offline data that allows us to overcome these challenges.
Additional Notation.
We use V to denote regularized design matrices given by $\mu\textbf{I}+\sum_{(x,a)}\phi(x,a)\phi(x,a)^{\top}$. We define $\textbf{D}_{n,i}=I-\mu\textbf{V}_{n,i}^{-1}$ and $\overline{\textbf{D}}_{N,i}=\frac{1}{n}\sum_{r=1}^{n}\textbf{D}_{n,i}$. Denote by $\hat{\beta}_{n,1},\hat{\beta}_{n,2}$ independent estimates of $\beta_{n}$ from $\tau_{n}$. Let $\textbf{M}_{n}=\frac{1}{2}(\hat{\bm{\beta}}_{n,1}\hat{\bm{\beta}}_{n,2}^{\top}+\hat{\bm{\beta}}_{n,2}\hat{\bm{\beta}}_{n,1}^{\top})$ and $\overline{\textbf{M}}_{N}\leftarrow\frac{1}{N}\sum_{n=1}^{N}\textbf{M}_{n}$.
3 Estimating Latent Subspaces Offline